Using LMV, prove that log(1+x) > 2x/(1+x) , if x>0


1
Expert's answer
2022-02-07T10:48:46-0500

Solution:

Let "f(x)=\log (1+x)"  in [0, x]

Since f(x) satisfies the condition of L.M.V. theorem in [0, x] , there exists "\theta(0<\theta<1)"  such that

"\begin{aligned}

&\frac{f(x)-f(0)}{x-0}=f^{\prime}(\theta x) \\

&\Rightarrow \quad \frac{\log (1+x)}{x}=\frac{1}{1+\theta x}

\end{aligned}"

Now, "\quad 0<\theta<1, x>0 \Rightarrow \theta x<x"

 "\Rightarrow \quad 1+\theta x<1+x 
\ \Rightarrow \quad \frac{1}{1+\theta x}>\frac{1}{1+x} 
\\\Rightarrow \quad \frac{x}{1+\theta x}>\frac{x}{1+x} \ ...(i)"

Again "0<\theta<1, x>0"

"\begin{array}{ll}

\Rightarrow & \theta x>0 \\

\Rightarrow & 1+\theta x>1 \\

\Rightarrow & \frac{1}{1+\theta x}<1 \\

\Rightarrow & \frac{x}{1+\theta x}<x\ ...(ii)

\end{array}"

From (i) and (ii), we get,

"\frac{x}{1+x}<\log (1+x)<x"